OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (jdoe0001):

89 boys and girls total b + g = 89 boys is twice of girls PLUS 8 b = 2g + 8 --------------------- b + g = 89 b = 2g + 8 there, solve by substitution

OpenStudy (whpalmer4):

To solve by substitution, \[b + g = 89\]\[b+g-b=89-b\]\[g=89-b\]Now take\[b=2g+8\]and rewrite it substituting \((89-b)\) wherever you have \(g\):\[b=2(89-b)+8\]Expand that and solve for \(b\)